2. If rolling two 6-sided dice, find the probability of obtaining a sum of 7.

Please Explain
Answer #1

Answer probability of getting a sum of 7 = 1/6

When two 6 sided dice or a dice are rolled the outcome is given in table :

2,2) |(23) (2, 2) (2,L) G.5 3, 3 1,2 2) (s 6,11 62) 61S

Total no of outcomes = 36

Let E​​​​​​1 = event of getting a sum of 7

n(E​​​​​​1)= (1,6),(2,5),(3,4),(4,3),(5,2),(6,1) = 6

P(E​​​​​​1​​​​​)= probability of getting a sum of 7

P(E​​​​​​1​​​​​) =number of outcomes /total number of outcomes

P(E​​​​​​1) =6/36 =1/6